OpenStudy (anonymous):

If A , B and C are any three set then prove that A-(b^c)= (A-B) U (A-C) ... ^= intersection

OpenStudy (helder_edwin):

\[ \large A-(B\cap C)=A\cap(B\cap C)^c \] \[ \large =A\cap(B^c\cup C^c)=(A\cap B^c)\cup(A\cap C^c) \] \[ \large =(A-B)\cup(A-C) \]

OpenStudy (anonymous):

then...

OpenStudy (helder_edwin):

then what?

OpenStudy (helder_edwin):

isn't this your question?

OpenStudy (anonymous):

thxxx i was just confused thxxxx

OpenStudy (helder_edwin):

u r welcome